OpenStudy (anonymous):

A kayaker paddles with a power output of 60.0 W to maintain a steady speed of 1.50 m/s. a)(a) Calculate the resistive force exerted by the water on the kayak.b)(b) If the kayaker increases her power output by a factor of 1.3, and the resistive force due to the water remains the same, by what factor does the kayaker's speed change?

OpenStudy (anonymous):

P=f/v so f=P/v so 60w/1.5s = 40N A) = 40N v = P/f so 78/40 = 1.9m/s B) 1.9m/s the change is by factor of 1.27. 1.5m/s * 1.27 = 1.9m/s